# Quantum plasmonics model of refractive index sensing using photon correlations

### v1

@@ V Sensing performance
In Fig. 6, the sensing performance of the scattered intensity at driving wavelengths within the plasmon spectrum (Figs. 6(a), (b), and (c)) is compared to the performance at wavelengths within the Fano spectrum (Figs. 6(d), (e), and (f)). The sensing performance was investigated using the refractive index range $n=1.3330-1.3334$ (corresponding to a typical range in a biosensing context [2]), where the mean photocount has a linear dependence on $n$, for each $\lambda_{D}$, as shown in Fig. S1 of the Supplementary Information. The intensity sensitivities, Fig. 6(a) and Fig. 6(d), have a similar dependence on the driving wavelengths — on the left-hand side of the plasmon resonance (Fano resonance), PL (FL) is the optimal driving wavelength, while PR (FR) is the optimal driving wavelength on the right-hand side of the plasmon resonance (Fano resonance). In each spectrum, the sensitivity is on the order of $\sim 10^{-4}$ RIU-1 and reaches a maximum value at PR (FR). As summarized in Table 2, there is no significant enhancement in the sensitivity since $\mathcal{E}_{S}\approx 1$ at the optimal driving wavelengths. Hence, the scattered intensity at the Fano spectrum is not necessarily more sensitive to changes in the refractive index than the scattered
@@
Figure 7: (Color online) (a) Intensity–intensity refractive index sensitivity, (b) standard deviation (noise) in the $g^{(2)}(0)$ signal for one second of measurements, and (c) sensing resolution (i.e., limit of detection) at driving wavelengths within the Fano spectrum, obtained within the refractive index range, $n=1.3330-1.3334$. The sensing resolution can be improved by taking measurements for more than one second.
intensity at the plasmon spectrum, given that the scattering peaks at the two regions lead to mean photocounts with the same order of magnitude.
@@
The dependence of the noise on driving wavelengths within the plasmon and Fano spectrum are shown in Fig. 6(b) and Fig. 6(e), respectively. In each spectrum, the standard deviation in the mean photocount, $\sigma_{m}$, is on the order of $\sim 10^{-8}$, and reaches a maximum value at PM (FM). Hence, in order to minimize the noise, the wavelengths, PM (FM), should not be used as driving wavelengths. As shown in Fig. 6(e), the scattered light within the Fano spectrum has the same first-order photon statistics as the scattered light within the plasmon spectrum, i.e., coherent light. This is because the low mean photocount within $\mathrm{T}_{int}$, $\langle\hat{m}\rangle\sim 10^{-4}$, renders the photon antibunching effect on $\Delta m$, and thus $\sigma_{m}$, insignificant within the Fano spectrum, i.e., $(g^{(2)}(0)-1)\langle\hat{m}\rangle\approx 0$ in Eq. (23) and Eq. (27a), so that the noise within the two spectra remains comparable. As shown in Fig. 6(c) and Fig. 6(f), the sensing resolution is not optimal near PM (FM) due to the combined effect of the low sensitivity and high noise. However, the resolutions are on the same order of magnitude, $\sim 10^{-5}$ RIU at the inflection points, resulting in no significant enhancement, since $\mathcal{E}_{\Delta n}\approx 1$, as shown in Table 2. Finally, we note that the sensing resolutions shown in Fig. 6(c) and 6(f) correspond to one second of measurements, and that both can be improved in principle by simply performing temporal averaging over a larger set of measurements. This, however, may lead to an increase in other sources of noise in the signal, such as technical noise in the setup (e.g. vibrations, equipment response etc.), which would impact the sensing resolution.
@@
Figure 8: Effects of the statistical property of the scattered light on the noise level assuming that photocounts in the range $0$ 1.0 and $g^{(2)}(0)\approx 0.17$ ($g^{(2)}(0)=1$) at the Fano resonance (plasmon resonance) are obtainable within the integration time, $\mathrm{T}_{int}$.

### v2

@@ V Sensing performance
In Fig. 6, the sensing performance of the scattered intensity at driving wavelengths within the plasmon spectrum (Figs. 6(a), (b), and (c)) is compared to the performance at wavelengths within the Fano spectrum (Figs. 6(d), (e), and (f)). The sensing performance was investigated using the refractive index range $n=1.3330-1.3334$ (corresponding to a typical range in a biosensing context [2]), where the mean photocount has a linear dependence on $n$, for each $\lambda_{D}$, as shown in Fig. S1 of the Supplemental Material [57].
The intensity sensitivities, Fig. 6(a) and Fig. 6(d), have a similar dependence on the driving wavelengths — on the left-hand side of the plasmon resonance (Fano resonance), PL (FL) is the optimal driving wavelength, while PR (FR) is the optimal driving wavelength on the right-hand side of the plasmon resonance (Fano resonance). The dips show that intensity sensing should not be carried out around the wavelength corresponding to the plasmon peak (and Fano peak), where the intensity has a smaller variation with the refractive index, i.e. the derivative of the intensity with respect to the refractive index is small at a fixed wavelength near the initial peak position. This is a common feature of intensity sensing at a resonance peak in an intensity spectrum — the intensity at a fixed wavelength near the initial peak does not vary significantly with refractive index, while the intensity on either side of the fixed wavelength varies more significantly (as the peak red or blue shifts).
In each spectrum, the sensitivity is on the order of $\sim 10^{-4}$ RIU-1 and reaches a maximum value at PR (FR). As summarized in Table 2, there is no significant enhancement in the sensitivity since $\mathcal{E}_{S}\approx 1$ at the optimal driving wavelengths. Hence, the scattered intensity at the Fano spectrum is not necessarily more sensitive to changes in the refractive index than the scattered intensity at the plasmon spectrum, given that the scattering peaks at the two regions lead to mean photocounts with the same order of magnitude.
@@
Figure 7: (Color online) (a) Intensity–intensity refractive index sensitivity, (b) standard deviation (noise) in the $g^{(2)}(0)$ signal for one second of measurements, and (c) sensing resolution (i.e., limit of detection) at driving wavelengths within the Fano spectrum (wavelengths around the Fano peak only), obtained within the refractive index range, $n=1.3330-1.3334$. The sensing resolution can be improved by taking measurements for more than one second.
Figure 8: Effects of the statistical property of the scattered light on the noise level assuming that photocounts in the range $0.0-1.0$ and $g^{(2)}(0)\approx 0.17$ ($g^{(2)}(0)=1$) at the Fano resonance (plasmon resonance) are obtainable within the integration time, $\mathrm{T}_{int}$.
@@
The dependence of the noise on driving wavelengths within the plasmon and Fano spectrum are shown in Fig. 6(b) and Fig. 6(e), respectively. In each spectrum, the standard deviation in the mean photocount, $\sigma_{m}$, is on the order of $\sim 10^{-8}$, and reaches a maximum value at PM (FM). Hence, in order to minimize the noise, the wavelengths, PM (FM), should not be used as driving wavelengths. As shown in Fig. 6(e), the scattered light within the Fano spectrum has the same first-order photon statistics as the scattered light within the plasmon spectrum, i.e., coherent light. This is because the low mean photocount within $\mathrm{T}_{int}$, $\langle\hat{m}\rangle\sim 10^{-4}$, renders the photon antibunching effect on $\Delta m$, and thus $\sigma_{m}$, insignificant within the Fano spectrum, i.e., $(g^{(2)}(0)-1)\langle\hat{m}\rangle\approx 0$ in Eq. (23) and Eq. (27a), so that the noise within the two spectra remains comparable. As shown in Fig. 6(c) and Fig. 6(f), the sensing resolution is not optimal near PM (FM) due to the combined effect of the low sensitivity and high noise. However, the resolutions are on the same order of magnitude, $\sim 10^{-5}$ RIU at the inflection points (which is within the range of those of start-of-the-art plasmonic sensors [8]), resulting in no significant enhancement, since $\mathcal{E}_{\Delta n}\approx 1$, as shown in Table 2. Finally, we note that the sensing resolutions shown in Fig. 6(c) and 6(f) correspond to one second of measurements, and that both can be improved in principle by simply performing temporal averaging over a larger set of measurements. This, however, may lead to an increase in other sources of noise in the signal, such as technical noise in the setup (e.g. vibrations, equipment response etc.), which would impact the sensing resolution.